# Scikit-fingerprints: easy and efficient computation of molecular fingerprints in Python

### v1

@@ Abstract
In this work, we present scikit-fingerprints, a Python package for computation of molecular fingerprints for applications in chemoinformatics. Our library offers an industry-standard scikit-learn interface, allowing intuitive usage and easy integration with machine learning pipelines. It is also highly optimized, featuring parallel computation that enables efficient processing of large molecular datasets. Currently, scikit-fingerprints stands as the most feature-rich library in the Python ecosystem, offering over 30 molecular fingerprints. Our library simplifies chemoinformatics tasks based on molecular fingerprints, including molecular property prediction and virtual screening. It is also flexible, highly efficient, and fully open source.
@@
C3 Permanent link to Reproducible Capsule
@@ 1 Motivation and significance
Molecules are the basic structures processed in computational chemistry. They are most commonly represented as molecular graphs, which need to be converted into multidimensional vectors for the majority of processing algorithms, most prominently for machine learning (ML) applications. This is typically done with molecular fingerprints, which are feature extraction algorithms encoding structural information about molecules as vectors [1]. They are ubiquitously used in chemoinformatics, e.g. for chemical space diversity measurement [2, 3, 4] and visualization [5, 6], clustering [7, 8, 9, 10], virtual screening [11, 12], molecular property prediction [13, 14, 15], and many more [16, 17, 18, 19, 20, 21]. These chemoinformatics tasks, often relying on machine learning methods, are important for many real-life applications, particularly in de novo drug design. For properly assessing the performance of predictive models, train-test splitting is crucial, and molecular fingerprints can also be used there [22, 23, 24, 25, 26]. The performance of fingerprint-based models remains very competitive, even compared to state-of-the-art graph neural networks (GNNs) [14].
@@
Selection of the optimal fingerprint representation for a given application is nontrivial, and typically requires computing many different fingerprints [14], and may also require tuning their hyperparameters [27, 28]. Using multiple fingerprints at once often improves results, e.g., via concatenation [15] or data fusion [29, 30]. Processing large molecular datasets necessitates efficient implementations that leverage modern multicore CPUs. Python, the most popular language in chemoinformatics today, includes the scikit-learn library [31], which has become the de facto standard tool for machine learning tasks. The library is renowned for its intuitive and widely adopted API [32].
@@
Popular open source tools for computing molecular fingerprints, such as Chemistry Development Kit (CDK) [33], OpenBabel [34] or RDKit [35], are written in Java or C++, and unfortunately only RDKit has an official Python wrapper. None of them are compatible with scikit-learn API, and they only support sequential computation. Each of these tools also supports only a limited number of fingerprints.
@@
Here, we present scikit-fingerprints, a new Python library for easy and efficient computation of molecular fingerprints. It is fully scikit-learn compatible, enabling easy integration into ML pipelines as a feature extractor for molecular data. It offers optimized parallel computation of fingerprints, enabling processing of large datasets and experiments with multiple algorithms, like data fusion. We implemented over 30 different fingerprints, making it the most feature-rich library in Python ecosystem for molecular fingerprinting. Those include ones based only on molecular graph topology (2D), as well as those utilizing graph conformational structure (3D, spatial). It is fully open source, publicly available on PyPI [36] and on GitHub at https://github.com/scikit-fingerprints/scikit-fingerprints.
@@ 2.1 Software architecture
scikit-fingerprints is a Python package for computing molecular fingerprints, and it is aimed at chemoinformatics and ML workflows. Its interface is fully compatible with scikit-learn API [32], ensured by proper inheritance from scikit-learn base classes and comprehensive tests.
The package structure is shown in Figure 1. All functionality is contained in the skfp package, allowing easy imports. Base classes are in skfp.bases package, and they can be used to extend the functionality with new or customized fingerprints. skfp.datasets has functions for loading popular datasets, for easy benchmarking. skfp.preprocessing contains classes for preprocessing molecules before computing fingerprints, as described in Section 2.2.1. Fingerprints are represented as classes in package skfp.fingerprints. Lastly, skfp.utils contains additional utility classes, such as input type validators.
@@ 2.2 Software functionalities
scikit-fingerprints user-facing functionalities can be broken into preprocessing and fingerprint calculation. It also supports loading popular datasets. In addition, in contrast to existing software, we support efficient parallelism, and implement multiple measures for ensuring high code quality and security.
@@ 2.2.1 Preprocessing
Fingerprints take RDKit Mol objects as input to the .transform() method. However, for convenience, all 2D-based fingerprints also take SMILES input, converting them internally. If done multiple times, this entails a small performance penalty, so scikit-fingerprints offers MolFromSmiles and MolToSmiles classes for easier conversions.
SMILES representation for a molecule is not unique, and does not convey all information. In particular, incorrect or very unlikely molecules can be written in SMILES form. MolFromSmiles by design performs only basic checks, to enable reading arbitrary data. For expanded checks, we implemented MolStandardizer class. Since there is no one-size-fits-all solution for molecular standardization, we use the most broadly used sanity checks, recommended by RDKit [37]. This helps to ensure high data quality at the beginning of the pipeline.
All fingerprints utilizing conformational (3D, spatial) information require Mol input, with conformers calculated using RDKit, with conf_id property set. Conformer generation can be troublesome, with multiple different algorithms and settings available. ConformerGenerator class in scikit-fingerprints greatly simplifies this process, offering reasonable defaults. It attempts to maximize efficiency for easy molecules and minimize failure chance for complex compounds, based on ETKDGv3 algorithm [38], known to give excellent results [39].
@@ 2.2.2 Fingerprints calculation
Different molecular fingerprints are represented as classes, all inheriting from BaseFingerprintTransformer, and further from BaseSubstructureFingerprint for substructure fingerprints like Klekota-Roth [40] (see Figure 2). They are used as stateless transformer objects in scikit-learn, and used mainly via .transform() method. It takes a list of SMILES strings or RDKit Mol objects, and outputs a dense NumPy array [41] or sparse SciPy array in CSR format [42]. Various options, such as vector length for hashed fingerprints (e.g. ECFP [43]), binary/count variant, dense/sparse output etc. are specified by constructor parameters. This ensures full composability with scikit-learn constructs like pipelines and feature unions.
We implement over 30 different fingerprints of various types, e.g. circular ECFP [43] and SECFP [44], path-based Atom Pair [45] and Topological Torsion [46], substructure-based MACCS [47] and Klekota-Roth [40], physicochemical descriptors like USRCAT [48] and Mordred [49], and more. We used efficient RDKit subroutines, written in C++, e.g. for matching SMARTS patterns. A full list of implemented fingerprints is available in the scikit-fingerprints online documentation.
@@ 2.2.3 Parallelism
Since molecules can be processed independently when computing fingerprints, the task is “embarrassingly parallel” [50]. This means that we can efficiently utilize all available cores. To minimize inter-process communication, by default, input molecules are split into as many chunks as there are physical cores available, and processed in parallel by Python workers. We utilize Joblib [51], with Loky executor, which uses memory mapping to efficiently pass the resulting arrays between processes. Furthermore, by using sparse arrays and smaller chunk sizes, users can minimize the memory utilization for large datasets and fingerprints that yield long output vectors [28].
Furthermore, we support distributed computing with Dask [52], used as Joblib executor. This way, scikit-fingerprints can take advantage of large high-performance computing (HPC) clusters. All that is required is a single parameter passed to Joblib configuration, to connect to the Dask cluster [53].
@@ 2.2.4 Datasets loading
Fingerprints are often used in the context of molecular property prediction on standardized benchmarks. In particular, they constitute strong baselines, often outperforming complex graph neural networks (GNNs) [13, 17, 14]. Therefore, their easy usage is important for fair evaluation of advancements in graph classification.
We utilized HuggingFace Hub [54, 55] to host datasets. It offers easy downloading, caching, and loading datasets, with automated compression to Parquet format. Currently, the most widely used MoleculeNet [56] benchmark has been integrated, and further datasets can be easily added with unified interface. Users can load datasets similarly to scikit-learn example datasets. For example, loading the BBBP dataset from MoleculeNet uses the function load_bbbp().
@@ 2.2.5 Code quality and CI/CD
We implemented a comprehensive suite of 196 unit and integration tests. They use PyTest framework [71], and are run automatically on GitHub Runners as a part of CI/CD process. Passing all tests is required to merge the code into the master branch. We run tests on a full matrix of operating systems (Linux, Windows, MacOS) and Python versions (from 3.9 to 3.12), ensuring proper execution in different environments.
@@ 3.1 Parallel computation
Computing fingerprints in parallel is useful for all molecular tasks, in particular for large databases in virtual screening. To illustrate the capability of scikit-fingerprints in this regard, we compute fingerprints for popular HIV dataset from MoleculeNet benchmark [56]. It contains a wide variety of molecules for a medicinal chemistry data, including organometallics, small and large molecules, some atoms with very high number of bonds etc. We limit the data to 10 thousand molecules, due to high computational time of running the benchmark multiple times for many data sizes and fingerprints. Code is available in the GitHub repository, in benchmarking directory.
@@
As an example, we present the timings for the PubChem fingerprint [72] in Figure 3, commonly used for virtual screening. Speedup for all fingerprints  [^1] We omit Pharmacophore fingerprint due to excessive computation time. Due to checking multiple SMARTS patterns for all atoms, it is by far the slowest fingerprint. is shown in Figure 4, when using 16 cores and 10 thousand molecules. Speedup is defined as a ratio of sequential to parallel computation time. We calculate those times as an average of 5 runs, using a machine with Intel Core i7-13700K 3.4 GHz CPU. For 3D fingerprints, we do not include the conformer generation time.
PubChem fingerprint clearly benefits from parallelism, with time decreasing with almost perfect speedup for more cores. This behavior is typical in particular for all substructure-based fingerprints, which have to check numerous SMARTS patterns for each molecule. This gain is especially significant for larger datasets.
High speedup values indicate that a significant majority of fingerprints benefits from parallelism, with Klekota-Roth achieving the greatest improvement. In general, computationally costly ones like SECFP or Mordred gain the most. Only the fastest ones like ECFP or Atom Pair have speedup less than 1, meaning slower computation than the sequential one. However, we did not tune the number of cores here, and using 2 or 4 could be more beneficial for those fingerprints given this amount of data.
@@ 3.2 Sparse matrix support
Molecular fingerprints are often extremely sparse, therefore using proper representation can result in large savings in memory usage, compared to dense arrays. Differences are particularly significant for large datasets, which are typical for virtual screening or similarity searching.
@@
Clearly, fingerprints greatly benefit from sparse representations, with density of arrays around just 1-2%. In particular, the very popular ECFP and FCFP fingerprints [43] are among those benefitting the most. The Klekota-Roth fingerprint [40], which is quite long for a substructure-based fingerprint, obtains a reduction from almost 2 GB RAM to just 23 MB, i.e. 88.2 times. Those savings would be even more important during hyperparameter tuning of downstream classifiers, when many copies of the data matrix are created. Using sparse representation did not negatively impact computation time, compared to the dense one.
@@ 3.3 Molecular property prediction
scikit-fingerprints can greatly simplify the process of classifying molecules. We show a part of a pipeline in Listing LABEL:code:molecular_property_prediction, responsible for computing ECFP fingerprints from SMILES strings and their classification. For brevity, we omit loading the data, which is just standard Pandas code for CSV files.
Inputs can be any sequences that consist of SMILES strings or RDKit Mol objects, e.g. Python lists, NumPy arrays, or Pandas series. Since ECFPFingerprint is a stateless transformer class, it uses an empty .fit() method in the pipeline. The code is also parallelized, requiring only the n_jobs parameter.
@@ 3.4 Fingerprint hyperparameter tuning
Most papers in the literature neglect hyperparameter tuning for molecular fingerprints, only tuning downstream classifiers. We conjecture that this is also due to the lack of easy to use and efficient software for computing fingerprints. The works that do perform such tuning [27, 28] indicate that it is indeed beneficial.
We performed hyperparameter tuning for all 2D fingerprints on MoleculeNet single-task classification datasets [56], using scaffold split provided by OGB [73]. Only the pharmacophore fingerprint was omitted due to excessive computation time for some molecules. A Random Forest classifier with default hyperparameters was used, in order to isolate the tuning improvements to just fingerprints. In Table 3, we report area under receiver operating characteristic curve (AUROC) values obtained when using tuned hyperparameters, improvement from tuning compared to the default parameters, and average gain over all datasets. Due to space limitations, we present the results for five fingerprints that had the highest average gain. They can be therefore considered the methods with the highest tunability [74]. Hyperparameter grids and code are available in the GitHub repository, in benchmarking directory.
Tuning fingerprints results in considerable gains, as high as 5.8% AUROC in case of RDKit fingerprint [75] on BBBP dataset. Notably, substructure-based Ghose-Crippen fingerprint [76] gains 4% AUROC on average, using feature counts instead of binary indicators. This signifies that further research in this area, utilizing scikit-fingerprints, would be highly beneficial.

### v3

@@ Abstract
In this work, we present scikit-fingerprints, a Python package for computation of molecular fingerprints for applications in chemoinformatics. Our library offers an industry-standard scikit-learn interface, allowing intuitive usage and easy integration with machine learning pipelines. It is also highly optimized, featuring parallel computation that enables efficient processing of large molecular datasets. Currently, scikit-fingerprints stands as the most feature-rich library in the open source Python ecosystem, offering over 30 molecular fingerprints. Our library simplifies chemoinformatics tasks based on molecular fingerprints, including molecular property prediction and virtual screening. It is also flexible, highly efficient, and fully open source.
@@
C3 Permanent link to Reproducible Capsule N/A
@@ 1 Motivation and significance
Molecules are the basic structures processed in computational chemistry. They are most commonly represented as molecular graphs, which need to be converted into multidimensional vectors for the majority of processing algorithms, most prominently for machine learning (ML) applications. This is typically done with molecular fingerprints, which are feature extraction algorithms that encode structural information about molecules as vectors [1]. They are widely used in chemoinformatics, e.g. for chemical space diversity measurement [2, 3, 4] and visualization [5, 6], clustering [7, 8, 9, 10], virtual screening [11, 12], molecular property prediction [13, 14, 15], and many more [16, 17, 18, 19, 20, 21]. These chemoinformatics tasks, which often rely on machine learning methods, are important for many real-life applications, particularly drug design. For properly assessing the performance of predictive models, train-test splitting is crucial, and molecular fingerprints can also be used there [22, 23, 24, 25, 26]. The performance of fingerprint-based models remains very competitive, even compared to state-of-the-art graph neural networks (GNNs) [14]. Hybrid molecular property prediction models are also a subject of recent research, combining molecular fingerprints with GNNs [27, 28, 29, 30], transformers [31, 32], or autoencoders [33].
@@
The selection of the optimal fingerprint representation for a given application is nontrivial. It typically requires the computation of many different fingerprints [14], and may also require tuning their hyperparameters [34, 35]. Using multiple fingerprints at once often improves results, e.g. via concatenation [15] or data fusion [36, 37]. Processing large molecular datasets necessitates efficient implementations that leverage modern multicore CPUs. Python, the most popular language in chemoinformatics today, includes the scikit-learn library [38], which has become the de facto standard tool for tabular machine learning tasks, and deep learning frameworks like PyTorch [39]. Scikit-learn in particular is renowned for its intuitive and widely adopted API [40].
@@
Popular open source tools for computing molecular fingerprints, such as Chemistry Development Kit (CDK) [41], Open Babel [42], and RDKit [43], are written in Java or C++. None of them are compatible with the scikit-learn API, and their Python wrappers can be cumbersome to work with. They also offer no or very limited support for parallel computation.
@@
Here, we present scikit-fingerprints, a new Python library for easy and efficient computation of molecular fingerprints. It is fully scikit-learn compatible, enabling easy integration into ML pipelines as a feature extractor for molecular data. It offers optimized parallel computation of fingerprints, enabling the processing of large datasets and experiments with multiple algorithms. We implemented over 30 different fingerprints, making it the most feature-rich library in the open source Python ecosystem for molecular fingerprinting. Those include those based only on molecular graph topology (2D), as well as those utilizing graph conformational structure (3D, spatial). It is fully open source, publicly available on PyPI [44] and on GitHub at https://github.com/scikit-fingerprints/scikit-fingerprints.
@@ 2.1 Software architecture
scikit-fingerprints is a Python package for computing molecular fingerprints, designed for chemoinformatics and ML workflows. Its interface is fully compatible with scikit-learn API [40], ensured by proper inheritance from scikit-learn base classes and comprehensive tests.
The package structure is shown in Figure 1. All functionality is contained in the skfp package, allowing easy imports. The base classes are in skfp.bases package, and they can be used to extend the functionality with new or customized fingerprints. skfp.datasets has functions to load popular datasets for easy benchmarking. skfp.preprocessing contains classes for preprocessing molecules before computing fingerprints, as described in Section 2.2.1. Fingerprints are represented as classes in package skfp.fingerprints. Lastly, skfp.utils contains additional utilities, such as input type validators.
@@ 2.2 Software functionalities
User-facing functionalities can be divided into preprocessing and fingerprint calculation. It also supports loading popular datasets. In addition, in contrast to existing software, we support efficient parallelism and implement multiple measures for ensuring high code quality and security.
@@ 2.2.1 Preprocessing
Fingerprints take RDKit Mol objects as input to the .transform() method. However, for convenience, all 2D-based fingerprints also take the SMILES input, converting them internally. If done multiple times, this entails a small performance penalty, so scikit-fingerprints offers MolFromSmiles and MolToSmiles classes for easier conversions.
SMILES representation for a molecule is not unique, and there are various non-standard extensions to this format [45, 46, 47]. In particular, incorrect or very unlikely molecules can be written in SMILES form. For example, string “[H]=[H]” is a syntactically correct SMILES, but is not a chemically valid molecule. MolFromSmiles by design performs only basic sanitization checks, to enable reading arbitrary data. For expanded checks, we implement the MolStandardizer class. Since there is no one-size-fits-all solution for molecular standardization, we use the most widely used standardization steps, recommended by RDKit [48]. This helps ensure high data quality at the beginning of the pipeline.
All fingerprints utilizing conformational (3D, spatial) information require Mol input, with conformers calculated using RDKit, with conf_id property set. Conformer generation can be troublesome, with multiple different algorithms and settings available. ConformerGenerator class in scikit-fingerprints greatly simplifies this process, offering reasonable defaults. It attempts to maximize efficiency for easy molecules and minimize the chance of failure for complex compounds, based on the ETKDGv3 algorithm [49], known to give excellent results [50].
@@ 2.2.2 Fingerprints calculation
Different molecular fingerprints are represented as classes, all inheriting from BaseFingerprintTransformer, and further from BaseSubstructureFingerprint for substructure fingerprints such as Klekota-Roth [51] (see Figure 2). They are used as stateless transformers in scikit-learn and used mainly via the .transform() method. It takes a list of SMILES strings or RDKit Mol objects, and outputs a dense NumPy array [52] or a sparse SciPy array in CSR format [53]. Various options, such as vector length for hashed fingerprints (e.g. ECFP [54]), binary/count variant, dense/sparse output etc. are specified by constructor parameters. This ensures full composability with scikit-learn constructs like pipelines and feature unions.
We implement more than 30 different fingerprints of various types, e.g. circular ECFP [54] and SECFP [55], path-based Atom Pair [56] and Topological Torsion [57], substructure-based MACCS [58] and Klekota-Roth [51], physicochemical descriptors such as EState [59] and Mordred [60], and more. We used efficient RDKit subroutines, written in C++, e.g. for matching SMARTS patterns. A complete list of implemented fingerprints is available in scikit-fingerprints online documentation.
@@ 2.2.3 Parallelism
Since molecules can be processed independently when computing fingerprints, the task is embarrassingly parallel [61]. This means that we can efficiently utilize all available CPU cores. To minimize inter-process communication, by default, input molecules are split into as many chunks as there are cores available, and processed in parallel by Python workers. We utilize Joblib [62], with the Loky executor, which uses memory mapping to efficiently pass the resulting arrays between processes. Furthermore, by using sparse arrays and smaller chunk sizes, users can minimize memory utilization for large datasets and fingerprints that yield long output vectors [35].
Furthermore, we support distributed computing with Dask [63], used as a Joblib executor. This way, scikit-fingerprints can take advantage of large high-performance computing (HPC) clusters. Connecting to the Dask cluster only requires setting a single parameter in the Joblib configuration [64].
@@ 2.2.4 Datasets loading
Fingerprints are often used in the context of molecular property prediction on standardized benchmarks. In particular, they constitute strong baselines, often outperforming complex graph neural networks (GNNs) [13, 17, 14]. Therefore, their easy usage is important for a fair evaluation of advances in graph classification.
We utilized HuggingFace Hub [65, 66] to host datasets. It offers easy downloading, caching, and loading datasets, with automated compression to Parquet format. Currently, the most widely used MoleculeNet [67] benchmark has been integrated, and additional datasets can be easily added with the unified interface. Users can load data sets as in scikit-learn. For example, loading the MoleculeNet BBBP dataset uses the function load_bbbp().
@@ 2.2.5 Code quality and CI/CD
We implemented a comprehensive suite of 196 integration and unit tests. They use the pytest framework [82], and are run automatically on GitHub Runners as a part of the CI/CD process. Passing all tests is required to merge the code into the master branch. We run tests on a full matrix of operating systems (Linux, Windows, macOS) and Python versions (from 3.9 to 3.12), ensuring proper execution in different environments.
@@ 3.1 Parallel computation
Since computing molecular fingerprints is an embarrassingly parallel task, it can very effectively utilize modern multicore CPU architectures, e.g. for large databases in molecular property prediction or virtual screening. To illustrate the capability of scikit-fingerprints in this regard, we calculate fingerprints for the popular HIV dataset from the MoleculeNet benchmark [67]. It contains a wide variety of molecules from medicinal chemistry, including organometallics, small and large molecules, some atoms with very high numbers of bonds, etc. For this experiment, we limit the data to 10 thousand molecules, due to the high computational time required to run the benchmark multiple times for many data sizes and fingerprints. The code is available in the GitHub repository, in benchmarking directory.
@@
As an example, we present the timings for the PubChem fingerprint [83], commonly used for virtual screening, in Figure 3. Speedup for all fingerprints  [^3] We omit Pharmacophore fingerprint due to excessive computation time. Due to the checking of multiple SMARTS patterns for all atoms, it is by far the slowest fingerprint. is shown in Figure 4, when using 16 cores and 10 thousand molecules. Speedup is defined as a ratio of sequential to parallel computation time. We calculate those times as an average of 5 runs, using a machine with Intel Core i7-13700K 3.4 GHz CPU. For 3D fingerprints, we do not include the conformer generation time.
PubChem fingerprint clearly benefits from parallelism, with time clearly decreasing when using more cores. This behavior is typical for more computationally heavy fingerprints, like substructure-based ones, which have to check numerous SMARTS patterns for each molecule. In particular, as visible in Figure 3, this gain appears for many data sizes. Even for just 2000 molecules, the time decreases from about 15 seconds to just about 2 seconds, which is much more convenient for interactive analyses and ad hoc queries, like searching for similar molecules.
High speedup values indicate that a significant majority of fingerprints benefit from parallelism, with Klekota-Roth achieving the greatest improvement. In general, computationally expensive ones like SECFP or Mordred gain the most. Only the fastest fingerprints, like ECFP or Atom Pair, have a speedup less than 1, meaning slower computation than the sequential one. However, we did not tune the number of cores, and using 6 or 8 could be enough for some fingerprints given this amount of data.
Lastly, in Figure 5 we provide a detailed speedup plot for six commonly used fingerprints of different types: hashed (ECFP [54] and RDKit [84]), substructural (MACCS [58] and PubChem [83]), and descriptors (EState [59] and Mordred [60]). Here, we could use the entire HIV dataset (about 41 thousand molecules), since the computational cost was much lower for only six fingerprints. Overall, fingerprints are well scalable with number of cores, particularly heavier ones like Mordred or PubChem. They achieve almost perfect linear speedup up to about 8 cores. Only the extremely fast ECFP fingerprint seems to be better suited for sequential computation, at least for a dataset of this size.
Figure 5: Speedup plot for selected fingerprints.
@@ 3.2 Sparse matrix support
Molecular fingerprints are often extremely sparse. Therefore, using proper representation can result in large memory savings, compared to dense arrays. Differences are particularly significant for large datasets, which are typical for virtual screening or similarity searching.
@@
Clearly, fingerprints greatly benefit from sparse representations, with a density of arrays around just 1-2%. In particular, popular ECFP and FCFP fingerprints [54] are among those that benefit the most. The Klekota-Roth fingerprint [51], which is quite long for a substructure-based fingerprint, obtains a reduction from almost 2 GB RAM to just 23 MB, i.e. 88.2 times. Those savings would be even more important during the hyperparameter tuning of downstream classifiers when many copies of the data matrix are created. Using a sparse representation did not negatively impact computation time, compared to the dense one.
@@ 3.3 Molecular property prediction
scikit-fingerprints can greatly simplify the process of classifying molecules. We show a part of a pipeline in Listing LABEL:code:molecular_property_prediction, responsible for computing ECFP fingerprints from SMILES strings and their classification. For brevity, we omit loading the data, which is just standard Pandas code.
Inputs can be any sequences that consist of SMILES strings or RDKit Mol objects, e.g. Python lists or Pandas series. Since ECFPFingerprint is a stateless transformer class, it uses an empty .fit() method in the pipeline. The code is also parallelized, requiring only the n_jobs parameter.
@@ 3.4 Fingerprint hyperparameter tuning
Most papers in the literature neglect hyperparameter tuning for molecular fingerprints, only tuning downstream classifiers. We conjecture that this is also due to the lack of easy-to-use and efficient software for computing fingerprints. The works that perform such tuning [34, 35] indicate that it is indeed beneficial.
We perform hyperparameter tuning for all 2D fingerprints on MoleculeNet single-task classification datasets [67], using the scaffold split provided by OGB [85]. Only the pharmacophore fingerprint was omitted due to the excessive computation time for some molecules. A Random Forest classifier with default hyperparameters was used, in order to isolate the tuning improvements to just fingerprints. In Table 3, we report the area under receiver operating characteristic curve (AUROC) values obtained when using tuned hyperparameters, improvement from tuning compared to the default parameters, and average gain over all datasets. Due to space limitations, we present the results for 5 fingerprints that had the highest average gain. They can therefore be considered as the methods with the highest tunability [86]. The hyperparameter grids and code are available in the GitHub repository, in benchmarking directory.
Tuning fingerprints results in considerable gains, as high as 5.8% AUROC in case of RDKit fingerprint [84] on BBBP dataset. Notably, substructure-based Ghose-Crippen fingerprint [87] gains 4% AUROC on average, using feature counts instead of binary indicators. This signifies that further research in this area, using scikit-fingerprints, would be highly beneficial.